If 3 more than a certain number is 2 less than a second number, the larger number minus the smaller number is equal to.
A)1
B)3
C)5
D)7

Answer by math_helper(2461)


C) 5.

To see this algebraically, notice that x + 3 = y - 2 —> y - x = 3 + 2